[figure: 9.7 cm]

MAGNETIC PROPERTIES OF Pr[Fe(CN)_{6}].5H_{2}O

A. Zentko^{a}, M. Bokor, K. Csach, V. Kavečanský^{a}, M. Lukáčová^{a}, M. Maryško^{b}, Z. Mitróová^{a}, M. Zentková^{a}

aInstitute of Experimental Physics, SAS, Watsonova 47, 043 53 Košice, Slovak Republic

bInstitute of Physics ASCR, Na Slovance 2 Prague, Czech Republic c Solid State Physics and Optics of HAS, Budapest, Hungary

Rare earth ferricyanide Pr[Fe(CN)_{6}].5H_{2}O was prepared by reacting of mixture PrCl_{3} solution with K_{3}[Fe(CN)_{6}] aqueous solution at room temperature. The prepared compound was characterized by thermogravimetric and differential thermal analysis. The hexagonal crystal structure (space group P 63/m) was refined from powder x-ray measurements by Rietveld method using GSAS software program package. The zero– field–cooled magnetization curves deviates from the field–cooled magnetization measurements M(T) below certain temperature T_{f} = 14.5 K, and shows a broad maximum at about 12 K. Such a behavior of M(T) is typical for the small particles system or for non–interacting magnetic clusters. From M(T) curves were determined the saturation magnetic moment μ_{s} = 2.1μ_{B}, which corresponds very well with theoretical one μ_{s}^{teor} = 2.2μ_{B}, and the ferrimagnetic Curie temperature T_{c} = 15.2 K. Magnetic susceptibility χ obeys Curie–Weiss law very well above T = 50 K with μ_{eff} = 3.8μ_{B} and Θ = -18.6 K. The M(H) curves show hysteresis behavior at T = 4.5 K. Low–temperature 1H- -spin–lattice relaxation time and ^{1}H–solid–echo of NMR spectra reflect changes in magnetic state of the sample. Both second moments (M_{2}) and full widths and half maximums (ν_{1/2}) have to much greater slopes at T < 30 K than at higher temperatures.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : A. Zentko

Address for correspondence : Institute of Experimental Physics, Slovak Academy of Sciences, Watsonova 47, 043 53 Košice, Slovak Republic

Email address : zentko@saske.sk